MICROWAVE SPECTROSCOPY OF THE HEAVY-ATOM CARBENE ANALOGS: HSiI and DSiI

\underline{LU KANG}, Department of Natural Sciences, Union College, Barbourville, KY 40906; MO-HAMMED A. GHARAIBEH, DENNIS J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506; STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459.

The pure rotational spectra of 6 silicon isotopologues, of HSiI and of DSiI have been recorded in natural abundance by pulsed-jet Fourier transform microwave (FTMW) spectroscopy. Neon was passed over dry ice cooled H_{3}SiI or D_{3}SiI and introduced into the pulsed valve of the FTMW spectrometer. The HSiI and the DSiI were produced in-situ with a 1000 V DC-discharge nozzle. Only a-type transitions can be observed from 6 - 26 GHz; K_{a} = 0 transitions for the HSiI and K_{a} = 0 and 1 transitions for the DSiI isotopologues. All observed transitions were assigned unambiguously. The molecular structure and chemical bonding of this heavy atom carbene will be discussed.